OpenStudy (anonymous):

Water in a canal, 30 dm wide and 12 dm deep, is flowing with velocity of 100 km per hour. How much area will it irrigate in 30 minutes, if 8 cm of standing water is required?

OpenStudy (anonymous):

30*12=360==> area

ganeshie8 (ganeshie8):

careful about the units, may be convert everything into cm

OpenStudy (anonymous):

50km will be irrigated in 30 minutes

OpenStudy (anonymous):

1dm=10cm

ganeshie8 (ganeshie8):

\[\large 300 \times 120 cm^2 \times \dfrac{100\times 1000\times 100 cm}{60 ~min} = 6\times 10^9 cm^3/min\]

OpenStudy (anonymous):

it is 30 minutes

OpenStudy (anonymous):

300*120/8=4500

ganeshie8 (ganeshie8):

In 30 minutes : \[\large 6\times 10^9 cm^3/min \times 30 min = 180\times 10^9 cm^3 \] water flows from the canal into the irrigation land

OpenStudy (anonymous):

how 100*1000*100cm?.100km=10000m

OpenStudy (anonymous):

10000m=1000000cm=10 00 00 00

ganeshie8 (ganeshie8):

100km 100 * 1000m 100 * 1000*100cm

ganeshie8 (ganeshie8):

i have changed everything into `cm`

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u r right:)

ganeshie8 (ganeshie8):

we're not done with the problem yet we got volume of water that went through the canal in 30 minutes, but we don't know how much land it will irrigate right ?

ganeshie8 (ganeshie8):

how much land can you irrigate with \(\large 180 \times 10^9 cm^3\) of water ?

ganeshie8 (ganeshie8):

an ideas ?

OpenStudy (anonymous):

180*10^9cm^3/8

ganeshie8 (ganeshie8):

again, pay attention to the units

ganeshie8 (ganeshie8):

180*10^9cm^3/8cm okay ?

OpenStudy (anonymous):

22.5*10^9*cm^3

ganeshie8 (ganeshie8):

units >.<

ganeshie8 (ganeshie8):

when u divide cm^3 by cm, what do u get ?

OpenStudy (anonymous):

cm^2

ganeshie8 (ganeshie8):

so its 22.5*10^9*cm^2 lets convert it to km^2

OpenStudy (anonymous):

22.5*10^9cm^2

ganeshie8 (ganeshie8):

or maybe m^2

OpenStudy (anonymous):

225000

OpenStudy (anonymous):

225000m^2

ganeshie8 (ganeshie8):

100cm = 1m (100cm)^2 = 1m^2 10^4cm^2 = 1m^2

ganeshie8 (ganeshie8):

22.5*10^9*cm^2 = 22.5*10^5*m^2

ganeshie8 (ganeshie8):

= 2.25*10^6*m^2

ganeshie8 (ganeshie8):

= 2.25 km^2

OpenStudy (anonymous):

yesssssssss........